OpenStudy (anonymous):

What is a polynomial function in standard form with zeros 1, 2, -2, and -3?

jimthompson5910 (jim_thompson5910):

If the zeros are 1, 2, -2, and -3, then x = 1, x = 2, x = -2, x = -3 x-1=0, x-2=0, x+2 = 0, x+3=0 (x-1)(x-2)(x+2)(x+3) = 0 I'll let you expand and finish up

OpenStudy (anonymous):

(x-1)(x-2)(x+2)(x+3)=0 By expanding you will get \[x ^{4}+2*x ^{3}-7*x ^{2}-8*x+12=0\]

OpenStudy (anonymous):

how did you get that though

jimthompson5910 (jim_thompson5910):

what do you get when you expand (x-1)(x-2)

OpenStudy (anonymous):

x^2-3x+2 ?

jimthompson5910 (jim_thompson5910):

how about (x+2)(x+3)

OpenStudy (anonymous):

x^2+5x+6

jimthompson5910 (jim_thompson5910):

good on both, so that means (x-1)(x-2)(x+2)(x+3) = 0 turns into (x^2-3x+2)(x^2+5x+6) = 0 then you expand that to get x^2(x^2+5x+6) - 3x(x^2+5x+6) + 2(x^2+5x+6) = 0 I'll let you do the next step

OpenStudy (anonymous):

x^4+5x^3-3x^3-15x^2-18x+2x^2+10x+12

OpenStudy (anonymous):

x^4+5x^3+6x^2-3x^3-15x^2-18x+2x^2+10x+12

jimthompson5910 (jim_thompson5910):

x^2(x^2+5x+6) - 3x(x^2+5x+6) + 2(x^2+5x+6) = 0 x^2*x^2+x^2*5x+x^2*6 - 3x*x^2- 3x*5x- 3x*6 + 2*x^2+2*5x+2*6 = 0 x^4+5x^3+6x^2 - 3x^3- 15x^2- 18x + 2x^2+10x+12 = 0 so far, so good, keep going

jimthompson5910 (jim_thompson5910):

now combine like terms and you're done

OpenStudy (anonymous):

x^4+2x^3-7x^2-8x+12

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

Yeah ! Keep going Little_Kay

OpenStudy (anonymous):

Thank you !

jimthompson5910 (jim_thompson5910):

you're welcome